OpenStudy (anonymous):

Name the function whose graph is the same as the graph of y=cosx with a phase shift of π/2.

OpenStudy (anonymous):

sin? not sure this is a guess hah

OpenStudy (anonymous):

cos not sin

OpenStudy (anonymous):

shift to right or left????

OpenStudy (anonymous):

no i meant if you were to phase shift by pi /2 it would be sin(x)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

either way the graph would be sin of x

OpenStudy (anonymous):

if it was shift to left or right

OpenStudy (anonymous):

I'm no idea , I cppy from book

OpenStudy (anonymous):

i assumed the answer BEFORE i went here lol

OpenStudy (ybarrap):

cos(x) = sin(x + pi/2) = cos(x)sin(pi/2) + sin(x)cos(pi/2) so Ans is sin(x+pi/2)

OpenStudy (anonymous):

i feel like it would be sinx also, because if you shift the graph pi/2 either way, it will end up being sinx

OpenStudy (anonymous):

i'm pretty sure it is sin(x) however i wouldn't just simply assume i'm sure there is work that you can do to probe it

OpenStudy (anonymous):

\[y=\sin x\]after phase shift\[y=\sin(\frac{\pi}{2} - x)=>y=cosx\]

OpenStudy (ybarrap):

note that sin(pi/2) = 1 and cos(pi/2) = 0 And sin(a+b) = sin(a)cos(b) + cos(a)sin(b)

OpenStudy (anonymous):

It should be -sin(x). cos(x) - pi/2 has us starting at y=0 but we're heading in the negative y direction which corresponds to -sin(x).

OpenStudy (anonymous):

thank all your guy

OpenStudy (ybarrap):

np